### v1

# HARMONIC: Cognitive and Control Collaboration in Human-Robotic Teams

Abstract
This paper presents a novel approach to multi-robot planning and collaboration. We demonstrate a cognitive strategy for robots in human-robot teams that incorporates metacognition, natural language communication, and explainability. The system is embodied using the HARMONIC architecture that flexibly integrates cognitive and control capabilities across the team. We evaluate our approach through simulation experiments involving a joint search task by a team of heterogeneous robots (a UGV and a drone) and a human. We detail the system’s handling of complex, real-world scenarios, effective action coordination between robots with different capabilities, and natural human-robot communication. This work demonstrates that the robots’ ability to reason about plans, goals, and attitudes, and to provide explanations for actions and decisions are essential prerequisites for realistic human-robot teaming.
I Introduction
In today’s rapidly evolving technological landscape, robots are becoming increasingly ubiquitous across various domains, from manufacturing and healthcare to search and rescue operations [1, 2]. This proliferation has given rise to a growing number of heterogeneous robot teams, where machines with diverse capabilities collaborate to achieve common goals. To harness these distributed and varied problem-solving capabilities, research in Multi-Agent Planning (MAP) has focused on developing sophisticated planning and collaboration strategies.
Despite significant advancements in multi-robot planning and collaboration [3, 4], a number of issues related to human-robotic teaming remain relatively unexplored [5]. For robots to serve as trusted partners in human-robotic teams, they must demonstrate a number of cognitive and metacognitive abilities, including:
  • developing and enhancing an understanding of team tasks and team organization;
  • developing and enhancing specific capabilities, involving the responsibilities, preferences, and actions of self and other team members [6];
  • maintaining and using an episodic memory of past activities, as when leveraging past experiences to make informed predictions about future needs and actions;
  • interpreting language, visual perception, and haptic perception in terms of an underlying formal model of the world (ontology); [7]
  • communicating with teammates using interpreted, meaningful natural language.
Figure 1: An overview of the framework, showing the Strategic and Tactical components representing the high-level planning and low-level execution, respectively.
The above capabilities will facilitate complex team interactions, attention management, goal-setting, planning (including incorporating a variety of MAP strategies), and overcoming unexpected challenges in a way that humans can readily interpret. The integration of verbal and visual communication enhances the overall effectiveness of skill transfer and knowledge sharing within human-robot teams. Furthermore, meaning-oriented natural language communication allows robots to provide causal explanations for their actions and decisions, thereby enhancing overall transparency and explainability, which are stepping stones to trust [7]. A growing number of robots today are based on foundational models, such as Large Language Models (LLMs) [8, 9] and Vision-Language-Action (VLA) models [10, 11]. These black-box technologies contribute to several components of our architecture. However, we make sure that the modules whose operation and output should be explanatory to humans will be fully glass-box inspectable and explainable in human terms both to team members and to external inspectors.
We present our approach to human-robot teaming using the example of a team comprised of two robots and a human executing a joint task. We demonstrate the robots’ knowledge of team organization as well as, their ability to (a) dynamically negotiate tasks and allocate responsibilities, (b) select and modify plans, and (c) adapt language communication to both the situation and the needs of their interlocutors. We implement these robotic capabilities using the novel HARMONIC architecture that operates in parallel at two levels – the strategic cognitive level, which supports operations requiring reasoning, and the tactical control level, which supports skill-based, ”automatic,” reflexive operation.
Specifically, in this paper, we report:
  • A cognitive strategy for multi-robot planning and execution with metacognition, communication, and explainability. This includes the robots’ natural language understanding and generation, their reasoning about plans, goals, and attitudes, and their ability to explain the reasons for their own and others’ actions.
  • The distributed HARMONIC architecture, which enables a flexible integration of cognitive and robotic control and execution capabilities.
  • Results of simulation experiments on a search task carried out by heterogeneous robots, which showcases the human-robot team’s ability to handle complex, real-world scenarios.
Section II briefly describes the state of the art. Sections III-V present the HARMONIC framework and the implementation of its strategic and tactical levels. Section VI describes the organization of planning tasks, and Section VII presents our initial evaluation of HARMONIC on a team search task.
II Background and Related Work
Multi-Agent Planning (MAP) in Heterogeneous Multi-Robot (HMR) teams encompasses a spectrum of strategies ranging from centralized to decentralized approaches, with varying degrees of inter-agent coupling. In centralized planning, a single agent orchestrates the entire team’s actions, potentially achieving globally optimal solutions but facing scalability challenges in larger teams. Conversely, decentralized methods enable autonomous plan negotiation among agents, offering enhanced scalability and adaptability at the risk of local suboptimality [12]. Brafman et al. [13] quantify this trade-off, noting that centralized planning in closely coupled multi-agent systems is exponentially more complex than solving independent problems in decoupled systems.
To balance these extremes, modern HMR teams often implement hybrid MAP strategies. These may include centralized hierarchical or collaborative planning coupled with decentralized execution [14], dynamic switching between local and global plans based on contextual factors [3], or centralized conflict resolution alongside decentralized planning [15]. Advanced implementations of these hybrid approaches include hierarchical planning frameworks, auction-based task allocation systems [16], and distributed partial-order planners [17], which have demonstrated significant collaborative capabilities in complex multi-robot scenarios.
Intra-team communication in an HMR team involves both inter-robot and human-robot interactions, each requiring distinct protocols. To date, most research has focused on inter-robot communication using standardized Agent Communication Languages to facilitate efficient information exchange [18]. The implementation of frameworks, such as CORA (Core Ontologies for Robotics and Automation) [19], while enhancing semantic interoperability and knowledge sharing among robotic agents, cannot support natural, human-level communication. To engender human trust, human-robot interactions must be carried out in everyday natural language.
The communication modalities currently employed in HMR teams, while effective for preprogrammed interactions between robots, Lemon et al. [20] and Natarajan et al. [5] support the need for a unified, human-centric approach to team communication. Rather than maintaining separate protocols for inter-robot and human-robot interactions, a common cognitive architecture that emulates human-like communication and reasoning offers a more cohesive solution, emulating human team dynamics.
Early work by Clark et al. [21] proposed a framework for this process, which was later adapted by Klein et al. [22] to team coordination, introducing the notion of common ground. These studies highlight essential capabilities for robotic team members, including maintaining shared understanding, negotiating goals and plans, managing attention, and handling commands and requests. Subsequent pioneering research in dialogue processing systems [23, 24, 25] and grounding [26, 27, 28, 29] further developed these ideas, laying the groundwork for addressing communication challenges in human-robot teams. Results of this research and our own prior work on team organization [30] and natural language communication [7] guided the implementation of cognitive processes for team organization and communication in HARMONIC.
III The HARMONIC Framework
HARMONIC (Human-AI Robotic Team Member Operating with Natural Intelligence and Communication), shown in Fig. 1, is a dual-control cognitive-robotic architecture that integrates strategic, cognitive-level decision-making with tactical, skill-level robot control. It builds upon and extends the concept of hybrid control systems and architectures, as discussed in a comprehensive review by Dennis et al. [31]. It also represents an advancement over the type 2 integration approach employed in the DIARC framework [32, 33]. Unlike DIARC, where the strategic layer is embedded as a subsystem within the tactical layer to enable concurrent and dynamic operations, HARMONIC introduces a more sophisticated integration of these components.
The strategic and tactical architectural layers of HARMONIC are connected by a bidirectional interface for seamless communication. The strategic layer, built upon the OntoAgent cognitive architecture [34, 35, 7], encompasses modules for attention management, perception interpretation, and decision-making (see Section. IV). It employs both utility-based and analogical reasoning, enhanced by metacognitive abilities. This layer prioritizes strategic goals, manages plan agendas, and selects actions while continuously monitoring their execution. It also facilitates team-oriented operations, including communicating in natural language and explaining decisions.
The tactical layer is responsible for robot control at a lower level, processing sensor inputs, and planning motor actions to execute high-level commands from the strategic layer. It employs controllers, algorithms, and models to translate abstract commands into concrete robot actions. For instance, a command PICK(key,at-position1) to ”pick up a key on the ground” initiates a series of precise operations, including object identification, position determination, trajectory computation, and actuation. This layer also manages reactive responses, such as collision avoidance, through dedicated controllers. In the tactical layer, HARMONIC employs Behavior Trees (BTs) [36], which provide modular and robust representations for skills and low-level plans (see Section. V).
The architecture’s design aligns with Kahneman’s dual-system approach [37], with the strategic layer implementing System 2 (slow, deliberative reasoning) and the tactical layer implementing System 1 (fast, intuitive responses). This structure allows for dynamic scheduling and adaptation, enabling real-time adjustments to priorities and actions. It effectively handles computational delays, contingencies, safety concerns, and resource optimization through a combination of low-level planning and reactive algorithms.
IV OntoAgent for Strategic Level Planning
OntoAgent is a content-centric cognitive architecture incorporated in HARMONIC. OntoAgent is designed to support the development of social intelligent agents through computational cognitive modeling [35, 7, 34]. This approach emphasizes the need to acquire, maintain, and dynamically expand large-scale knowledge bases, which are essential for an agent’s perception interpretation, reasoning, and action. The architecture’s memory structure is divided into three main components: a Situation Model (SM) that contains currently active concept instances; a Long-Term Semantic Memory (LTS) that stores knowledge about instances of events and objects, and an Episodic Memory (LTE) that stores knowledge about instances of events and objects. OntoAgent supports goal-directed behavior through a goal agenda and a prioritizer that selects which goals to pursue. It typically uses stored plans associated with goals but can also engage in reasoning from first principles when necessary.
The architecture is built on a service-based infrastructure, with key services including perception interpretation, attention management, reasoning, and action rendering. OntoAgent can integrate both native services and imported external capabilities, such as robotic vision or test-to-speech modules. This flexibility allows it to support multiple perception channels, including language understanding, visual perception, and simulated interoception (Chapter 8 of [35]).
A crucial component of OntoAgent is OntoGraph, a knowledge base API that provides a unified format for representing and accessing knowledge across the system. OntoGraph supports inheritance, flexible organization of knowledge into ”spaces,” and efficient querying and retrieval. It implements a graph database view of knowledge, allowing for complex relational queries and supporting the representation of inheritance hierarchies.
OntoAgent places a strong emphasis on natural language understanding and meaning-based text generation. These capabilities rely on a semantic lexicon that links words and phrases with meanings grounded in a resident ontology. The language analyzer treats a large range of complex linguistic phenomena (lexical disambiguation, reference resolution, ellipsis reconstruction, new-word learning, etc. [7]) and generates ontologically-grounded interpretations of text meaning. The language generator, for its part, generates natural language utterances from ontologically grounded meaning representations that the agent creates as part of its reasoning about action. OntoAgent also supports ontological interpretation of visual percepts using an opticon whose entries link images to ontological objects.
The architecture supports goal-directed behavior through a goal agenda and a prioritizer that selects which goals to pursue. It typically uses stored plans associated with goals but can also engage in reasoning from first principles when necessary.
V Behavior Trees for Tactical Execution
Behavior Trees (BTs), originally developed for controlling non-player characters in video games, have found significant applications in robotics and AI [36]. BTs are directed trees that provide a flexible and intuitive way to design control actions and define task-planning hierarchies for agents. The flexibility and modularity of BTs make them particularly suitable for complex, hierarchical task planning in robotics and AI applications, allowing for easy modification and scaling of agent behaviors.
BTs in the tactical layer of HARMONIC enable effective reactive control and support the system’s safety and operational requirements in dynamic environments. The tactical layer also incorporates advanced control techniques such as whole-body compliant control [38, 36] with motion planning, path planning [39] etc. In addition to these, BTs can also be used for representing skills and low-level plans that can directly be shared between the robots[40, 41].
BTs typically use a blackboard architecture to maintain and access frequently checked condition variables [42, 43]. In the context of HARMONIC, a state manager keeps track of these variables on the tactical layer and allows for efficient querying and updating of the system’s state during operation based on the sensory inputs and action commands received from the strategic layer.
Figure 2: (a) Root, control flow, and execution nodes in BTs. (b) BT design template for the tactical layer of robots running the HARMONIC architecture. (c) Sample BT on UGV.
BTs for HARMONIC robots follow a design template prioritizing safety and needs [40], as shown in Fig. 2. Collision avoidance subtrees are placed leftmost, followed by robot needs subtrees, then action command subtrees triggered by the strategic layer. Fallback subtrees on the far right provide actions when the robot is idle or other conditions aren’t met, such as random walks or waiting at the base station. This structure leverages the left-to-right execution of BTs to ensure proper prioritization of tasks.
The translation of the action commands to the appropriate state flags and control variables is done through the action data communication APIs, which modify the placeholder variables in the State Manager on the tactical layer.
VI Development of Distributed Plans
In this section, we show how robots implemented in HARMONIC carry out a search task. They decide to act based on reasoning carried out in the strategic layer. Their plans (which are instances of ontological scripts) are divided into steps that are communicated to the tactical layer by means of action commands.
VI-A High-level planning
The ontologies of agents in HARMONIC contain scripts for complex events that a particular agent can perform or understand other agents performing. Plans are as instances of scripts, with parameter values set for the situation at hand. Plans are instantiated by executing one or more of ontologically stored meta-plans, that is, plans that HARMONIC robots instantiate to guide the generation of plans in their operational domain. One such meta-plan is the collaborative activity plan, whose purpose is to help agents operating in teams organize themselves to accomplish a goal.
The details of the collaborative activity plan vary based on the agent’s role in the team. For example, the team (or task) leader’s plan is typically more complex than the plans of its subordinates since team leaders select the plan to accomplish the target goal, resolve preconditions, and instruct their subordinates. Subordinates’ plans, by contrast, include awaiting instructions from the leader and, passively absorbing any other related information that comes their way.
The leader’s script for collaboration is shown below. Since this is a meta-plan, it can be applied to many specific plans and, starts with few details populated. Part of its task is to fill in task-specific details. Select comments are provided for this human-readable formalism. For further information, please refer to [34, 35, 7].
{codeblock}
@COLLABORATIVE-ACTIVITY (leader) [INIT] *identify-team-members [SELECT-PLAN] // Select a plan from available options. RUN *identify-candidate-plans RUN *select-plan [PRECONDITIONS] // Preconditions will be determined once // a plan has been picked. [SUGGEST-PLAN] // Propose the selected plan to the team. RUN NEW @PROPOSE-PLAN [RUN-PLAN] // The leader hasn’t picked a plan yet.
The subordinate’s script for collaboration is:
{codeblock}
@COLLABORATIVE-ACTIVITY (subordinate) [INIT] *identify-team-members [WAIT-FOR-PLAN] // Wait for a plan from the leader before // continuing. AWAIT \$.HAS-COLLABORATIVE-PLAN ISA @EVENT [RUN-PLAN] // The subordinate has no plan yet.
As the leader begins to process the meta-plan, it will select the domain plan to pursue. In the example below, the selected plan is SEARCH-FOR-LOST-OBJECT. The details of the selected plan can be placed into the [RUN-PLAN] section of collaborative activity, and any preconditions that must be resolved can be added to the [PRECONDITIONS] section. An excerpt from the leader’s script for COLLABORATIVE-ACTIVITY is shown below.
{codeblock}
@COLLABORATIVE-ACTIVITY (leader) [INIT] *identify-team-members [SELECT-PLAN] // Select a plan from available options. RUN *identify-candidate-plans RUN *select-plan [PRECONDITIONS] // Preconditions from the selected plan // to be checked. RUN NEW @REQUEST-OBJECT-TYPE RUN NEW @REQUEST-OBJECT-FEATURES RUN NEW @REQUEST-LAST-SEEN-AT RUN NEW @REQUEST-LOCATION-CONSTRAINED [SUGGEST-PLAN] // Propose the selected plan to the team. RUN NEW @PROPOSE-PLAN [RUN-PLAN] // Execute the selected plan, step by step. RUN NEW @SEARCH-FOR-LOST-OBJECT @SEARCH-FOR-LOST-OBJECT [SEARCH-ZONES] // Search in each ZONE the agent knows. // Stop searching when the object has been located. // Report the results (if necessary) to the team. FOR #ZONE-1 IN #LOCATION-1.SEARCHABLE-ZONE RUN ASYNC AWAIT *search INTERRUPT WHEN #OBJECT-1.LOCATION KNOWN RUN *consider-reporting
Once the leader has worked through the preconditions, it can move to SUGGEST-PLAN, in which it will share the domain plan with subordinates through natural language dialog. These agents will receive this information and adopt their assignment(s) within the suggested plan. This process results in the team adopting a shared collaborative plan.
VII Evaluation
VII-A Simulation Environment and Task Design
We evaluated the HARMONIC framework’s teaming strategy using a simulated search and retrieval task set in an apartment environment, as illustrated in Figure 5. The scenario involves a HMR team consisting of an unmanned ground vehicle (UGV), a drone, and a human team member serving as the team leader. The objective is to find a set of lost keys. The human initiates the task by remotely communicating with the robotic team members.
The simulated apartment contains zones with varied robot accessibility, necessitating coordinated efforts. We categorize these as: a) spaces accessible to both robots, b) elevated surfaces accessible to only the drones, and c) spaces below surfaces accessible only to the UGVs, as shown in Fig. 3. Each zone comprises waypoints - strategically defined locations that are dynamically utilized by the strategic layer. For this evaluation, we manually designated these waypoints and zones. However, in the future this process will be automated using machine learning models to classify and map zones based on spatial characteristics and accessibility, which will enhance the system’s adaptability to new environments.
Figure 3: Zones in the search environment
The simulation environment, developed using the Unity game engine, incorporates abstracted versions of robot behaviors to reduce complexity. This approach allowed us to focus on key aspects of the HARMONIC framework’s performance while simplifying the implementation of complex robot actions that are not central to our current work.
VII-B Heterogenous Robot Team Configuration
The simulations involved two robots, an Unmanned Ground Vehicle (UGV) and a drone (Fig. 4). These robots are both functionally and structurally heterogeneous [44], thus providing a diverse range of complementary perspectives and capabilities to the team. Both robots are equipped with a suite of sensors that enable them with localization, collision detection, and object recognition capabilities.
Figure 4: (a) UGV for ground-level exploration. (b) Parrot drone conducting an aerial scan of objects.
Each robot runs its own instances of the tactical and strategic modules, enabling decentralized planning and control within the team. For the current scenario, the UGV is the robot team leader at the task level, making it responsible for maintaining interactions with the human and the drone. Currently, the leader assignment is randomized. In future work, we intend to develop and evaluate heuristic methods for implementing dynamic team hierarchies. This approach will optimize leadership selection based on contextual factors and individual agent capabilities depending on the task.
In the tactical layer of both robots, sensor data is encoded into data frames and periodically shared with the strategic layer. Additionally, each robot incorporates reactive control strategies for collision avoidance, that are embedded within behavior trees in their respective tactical layers.
The UGV is a custom-designed robot featuring a manipulator mounted atop a multidirectional mobile platform, as illustrated in Fig. 4.a. The robot’s design allows it to access all ground-level locations and maneuver under certain objects, enhancing its capability to perform intricate search and manipulation tasks in confined spaces. Figure 2.c presents an excerpt from the Behavior Tree (BT) running on the Unmanned Ground Vehicle (UGV). This BT is based on the template shown in Figure 2.a.
Figure 5: (1) Chat transcript between Danny and the robots. (2) UGV’s (leader) Text Meaning Representations (TMRs) (3, 6) A sample Vision Meaning Representations (VMRs) of detected objects by the robots (5) Leader’s agenda. (4, 7) Complete thought transcripts of UGV and Drone.
The drone, on the other hand, can effectively perform search operations by scanning through a wide range of environments with sensors to scan the objects below it, as shown in Fig. 4.b. Its vertical mobility allows it to navigate over obstacles and access areas that may be challenging for the UGV (type-b zones).
Since the robots communicate in natural language, their dialog and actions are comprehensible to humans, and humans can intervene when necessary.
VII-C Collaborative Search for Lost Keys
The simulation environment is connected to a chatroom interface through which the human (Danny) communicates with the robots. A user interface (see Fig. 5) includes not only a chat window but additional widgets that display the robots’ interpretations of language (TMRs) and visual inputs(VMRs), natural language renderings of the robots’ thoughts, and their goal and plan agendas. This arrangement illustrates the transparency of the robots’ operations and provides insights into their cognitive processes. The discussion below makes reference to the labels in Fig. 5  [^1] A video of the complete simulation can be found at https://youtu.be/LpuE0yN0SEI
VII-C1 Task Initiation
The scenario begins with Danny, initiating the task by sending the robots a message (M1). This input triggers the team leader (UGV) to place a COLLABORATIVE-ACTIVITY on its agenda (Fig. 5.5) and launch the SEARCH-FOR-LOST-OBJECT plan. Concurrently, the drone awaits instructions.
VII-C2 Information Gathering
The UGV proceeds to verify preconditions for the SEARCH-FOR-LOST-OBJECT plan (Fig. 5.5). Although the object type is already known, information about its features and last-seen location would be useful. The UGV queries Danny for this information (M2/M4), and receives responses (M3/M5). The latter response prompts the system to prioritize the entry-way sub-zone (type-a zone) in the search sequence.
VII-C3 Search Execution
With preconditions met, the UGV instructs the drone to initiate the search (M6), and both robots begin exploring their assigned areas. The individual tactical modules on the robots control the search process using a waypoint strategy, while the strategic (cognitive) module maintains awareness of area existence without directly guiding robot navigation. This approach allows for efficient local path planning while preserving high-level planning in the strategic layer.
VII-C4 Communication and Coordination
Throughout the search, robots report their findings to each other and the human (M7-9). When a robot fails to locate the keys in a searched area, it communicates this to its partner.
VII-C5 Task Completion
The VMRs widget (Fig. 5.3 and Fig. 5.6) displays the object detection results that the strategic layer processes from sensing frames communicated by the tactical layer. During the search execution, the strategic module continuously analyzes the sensor data frames, grounding these VMRs against the instance of the KEY object stored in its episodic memory. When the features match, the search is halted by either of the robots (leader in this case), informs the team (M8), and reports to Danny (M9). Notably, the UGV uses different language constructs when communicating with the drone versus Danny in (M9), demonstrating the cognitive agent’s ability to generate context-appropriate language.
VIII Conclusion and Future Work
This work presents a step towards the development of cognitive robots capable of effectively collaborating in heterogeneous human-robot teams. The HARMONIC architecture, with its dual-layer approach that combines strategic cognitive processing and tactical control enables the embodiment of cognitive architectures enhancing robots to exhibit metacognitive abilities, including team task understanding, episodic memory utilization, and multimodal perception interpretation. By integrating natural language communication and reasoning capabilities, our approach facilitates complex team interactions, enhances explainability, and builds trust the key factors in human-robot collaboration.
The glass-box nature of our system, unlike black-box models based on LLMs or VLAs, provides transparency and interpretability to modules whose operation and output should be explainable to humans. Our demonstration system of a simulated search task involving two robots and a human teammate illustrates the potential for HARMONIC to be applied to real-world applications. Future research will focus on expanding the range of tasks and scenarios in which HARMONIC agents can successfully operate; enhancing the robots’ learning capabilities; conducting real-world trials using physical robots to validate the framework’s effectiveness in diverse operational contexts; developing collaborative online planning strategies; and, enhancing real-time adaptability in dynamic environments.
ACKNOWLEDGMENT
This research was supported in part by grant ##N00014-23-1-2060 from the U.S. Office of Naval Research. Any opinions or findings expressed in this material are those of the authors and do not necessarily reflect the views of the Office of Naval Research.

### v2

@@
 Abstract
-This paper presents a novel approach to multi-robot planning and collaboration.
-We demonstrate a cognitive strategy for robots in human-robot teams that incorporates metacognition, natural language communication, and explainability.
-The system is embodied using the HARMONIC architecture that flexibly integrates cognitive and control capabilities across the team.
-We evaluate our approach through simulation experiments involving a joint search task by a team of heterogeneous robots (a UGV and a drone) and a human.
-We detail the system’s handling of complex, real-world scenarios, effective action coordination between robots with different capabilities, and natural human-robot communication.
-This work demonstrates that the robots’ ability to reason about plans, goals, and attitudes, and to provide explanations for actions and decisions are essential prerequisites for realistic human-robot teaming.
+This paper introduces HARMONIC, a cognitive-robotic architecture that integrates the OntoAgent cognitive framework with general-purpose robot control systems applied to human-robot teaming (HRT).
+We also present a cognitive strategy for robots that incorporates metacognition, natural language communication, and explainability capabilities required for collaborative partnerships in HRT.
+Through simulation experiments involving a joint search task performed by a heterogeneous team of a UGV, a drone, and a human operator, we demonstrate the system’s ability to coordinate actions between robots with heterogeneous capabilities, adapt to complex scenarios, and facilitate natural human-robot communication.
+Evaluation results show that robots using the OntoAgent architecture within the HARMONIC framework can reason about plans, goals, and team member attitudes while providing clear explanations for their decisions, which are essential prerequisites for realistic human-robot teaming.
 I Introduction
-In today’s rapidly evolving technological landscape, robots are becoming increasingly ubiquitous across various domains, from manufacturing and healthcare to search and rescue operations [1, 2].
-This proliferation has given rise to a growing number of heterogeneous robot teams, where machines with diverse capabilities collaborate to achieve common goals.
-To harness these distributed and varied problem-solving capabilities, research in Multi-Agent Planning (MAP) has focused on developing sophisticated planning and collaboration strategies.
-Despite significant advancements in multi-robot planning and collaboration [3, 4], a number of issues related to human-robotic teaming remain relatively unexplored [5].
-For robots to serve as trusted partners in human-robotic teams, they must demonstrate a number of cognitive and metacognitive abilities, including:
+The proliferation of robots across various domains has given rise to a growing number of human-robot teams, where robots with various capabilities are employed by human operators to achieve complex mission goals [1].
+Current state-of-the-art robotic systems demonstrate remarkable tactical proficiency, from the sub-millimeter precision of surgical robots performing intricate surgical procedures [2] to the dynamic multi-terrain locomotion of quadrupedal [3] and humanoid platforms exhibiting sophisticated whole-body coordination [4].
+These systems also excel at integrating multi-modal sensory information through sensor fusion, simultaneously localizing themselves within complex environments, recognizing objects, tracking moving targets with increasing reliability, in adaptive policy learning and multi-robot planning and collaboration [5, 6].
+Despite these significant advancements in tactical execution, most state-of-the-art robots are limited to short-horizon planning and execution and lack the strategic capabilities necessary for long-horizon planning, reasoning, and true collaborative partnership with humans in complex decision-making scenarios [7].
+Importantly, effective human-robot teaming requires more than just tactical proficiency; to function as trusted partners within human-robot teams, they must exhibit a range of cognitive and metacognitive capabilities, including:
   • developing and enhancing an understanding of team tasks and team organization;
@@
   • communicating with teammates using interpreted, meaningful natural language.
-Figure 1: An overview of the framework, showing the Strategic and Tactical components representing the high-level planning and low-level execution, respectively.
-The above capabilities will facilitate complex team interactions, attention management, goal-setting, planning (including incorporating a variety of MAP strategies), and overcoming unexpected challenges in a way that humans can readily interpret.
-The integration of verbal and visual communication enhances the overall effectiveness of skill transfer and knowledge sharing within human-robot teams.
-Furthermore, meaning-oriented natural language communication allows robots to provide causal explanations for their actions and decisions, thereby enhancing overall transparency and explainability, which are stepping stones to trust [7].
-A growing number of robots today are based on foundational models, such as Large Language Models (LLMs) [8, 9] and Vision-Language-Action (VLA) models [10, 11].
-These black-box technologies contribute to several components of our architecture.
-However, we make sure that the modules whose operation and output should be explanatory to humans will be fully glass-box inspectable and explainable in human terms both to team members and to external inspectors.
+Figure 1: An overview of the framework, showing the Strategic and Tactical components representing the high-level planning (System 2) and low-level execution (System 1), respectively.
+Long-horizon functioning is within the scope of the field of cognitive architectures where much progress has been made in supporting complex agent team interactions, managing attention, setting goals, and conducting sophisticated planning, often incorporating diverse multi-agent planning strategies.
+Still, while cognitive architectures address key limitations in current robotic systems, they do not traditionally concentrate on the dynamic world of embodied agents.
+This has resulted in a notable disconnect between cognitive architectures and practical robotic control systems.
+Bridging this divide leverages the strengths of both these areas of research and development.
+The HARMONIC framework presented in this paper directly addresses this challenge through the use of a dual-layer architecture that combines strategic cognitive capabilities with tactical robotic control.
+This integration not only enables long-horizon planning but also enhances the scope of the robot’s reasoning capabilities, adds metacognitive awareness and social intelligence, which in turn supports human-level explanation that serves as a foundation for robots to function as trusted teammates [10], rather than mere tools, within human-robot teams.
+The cognitive architecture that we incorporate into HARMONIC as its strategic layer is OntoAgent [11, 12, 9].
+It is responsible for the aforementioned cognitive architecture capabilities, thus facilitating complex team interactions, attention management, goal-setting, sophisticated planning (including the incorporation of diverse multi-agent planning strategies), and addressing unexpected challenges in ways that humans can readily interpret.
+Additionally, OntoAgent integrates both verbal and visual communication, significantly enhancing skill transfer and knowledge sharing within human-robot teams, grounding, and supporting meaning-oriented natural language communication, enabling clear causal explanations for the robots’ actions and decisions.
+These capabilities greatly enhance transparency and explainability [10], establishing crucial foundations for trust in human-robot collaborations [9].
 We present our approach to human-robot teaming using the example of a team comprised of two robots and a human executing a joint task.
 We demonstrate the robots’ knowledge of team organization as well as, their ability to (a) dynamically negotiate tasks and allocate responsibilities, (b) select and modify plans, and (c) adapt language communication to both the situation and the needs of their interlocutors.
-We implement these robotic capabilities using the novel HARMONIC architecture that operates in parallel at two levels – the strategic cognitive level, which supports operations requiring reasoning, and the tactical control level, which supports skill-based, ”automatic,” reflexive operation.
-Specifically, in this paper, we report:
+We implement these robotic capabilities using the novel HARMONIC architecture that operates in parallel at two levels – the strategic cognitive level, which supports operations requiring reasoning, and the tactical control level, which supports skill-based, ”automatic,” reflexive operation of robots.
+Specifically, the key contributions of this paper are:
+  • We introduce HARMONIC, a cognitive-robotic architecture, which enables a flexible integration of OntoAgent [11], a cognitive architecture with low-level robot planning and control using Behavior Trees [13].
   • A cognitive strategy for multi-robot planning and execution with metacognition, communication, and explainability.
 This includes the robots’ natural language understanding and generation, their reasoning about plans, goals, and attitudes, and their ability to explain the reasons for their own and others’ actions.
-  • The distributed HARMONIC architecture, which enables a flexible integration of cognitive and robotic control and execution capabilities.
-  • Results of simulation experiments on a search task carried out by heterogeneous robots, which showcases the human-robot team’s ability to handle complex, real-world scenarios.
-Section II briefly describes the state of the art.
-Sections III-V present the HARMONIC framework and the implementation of its strategic and tactical levels.
-Section VI describes the organization of planning tasks, and Section VII presents our initial evaluation of HARMONIC on a team search task.
+  • Results of simulation experiments on a search task carried out by a robotic team, which showcases the human-robot team’s interactions and ability to handle complex, real-world scenarios.
+Section II briefly describes the state of the art and background.
+Section III introduces the HARMONIC framework and the implementation of its strategic and tactical levels.
+Section IV-B describes the task, the team configuration, and the development of distributed plans, and Section IV presents the evaluation.
 II Background and Related Work
-Multi-Agent Planning (MAP) in Heterogeneous Multi-Robot (HMR) teams encompasses a spectrum of strategies ranging from centralized to decentralized approaches, with varying degrees of inter-agent coupling.
-In centralized planning, a single agent orchestrates the entire team’s actions, potentially achieving globally optimal solutions but facing scalability challenges in larger teams.
-Conversely, decentralized methods enable autonomous plan negotiation among agents, offering enhanced scalability and adaptability at the risk of local suboptimality [12].
-Brafman et al.
-[13] quantify this trade-off, noting that centralized planning in closely coupled multi-agent systems is exponentially more complex than solving independent problems in decoupled systems.
-To balance these extremes, modern HMR teams often implement hybrid MAP strategies.
-These may include centralized hierarchical or collaborative planning coupled with decentralized execution [14], dynamic switching between local and global plans based on contextual factors [3], or centralized conflict resolution alongside decentralized planning [15].
-Advanced implementations of these hybrid approaches include hierarchical planning frameworks, auction-based task allocation systems [16], and distributed partial-order planners [17], which have demonstrated significant collaborative capabilities in complex multi-robot scenarios.
-Intra-team communication in an HMR team involves both inter-robot and human-robot interactions, each requiring distinct protocols.
-To date, most research has focused on inter-robot communication using standardized Agent Communication Languages to facilitate efficient information exchange [18].
-The implementation of frameworks, such as CORA (Core Ontologies for Robotics and Automation) [19], while enhancing semantic interoperability and knowledge sharing among robotic agents, cannot support natural, human-level communication.
-To engender human trust, human-robot interactions must be carried out in everyday natural language.
-The communication modalities currently employed in HMR teams, while effective for preprogrammed interactions between robots, Lemon et al.
+II-A Cognitive-Robotic Architectures
+There are relatively few cognitive architectures explicitly designed or effectively adapted for robotic applications due to the significant challenges involved in combining cognitive-level decision-making with motor control.
+Among these architectures, Soar [14], ACT-R [15], and DIARC [16] are prominent, each contributing distinct methods and philosophies toward cognitive autonomy in robotics.
+The Soar architecture provides symbolic cognitive processing through rule-based decision-making, facilitating general problem-solving and reasoning [14].
+It excels at clear, structured decision-making but requires extensive manual integration for perception and motor control, making it challenging to deploy in unpredictable or sensor-rich environments without substantial customization.
+The ACT-R architecture [15], on the other hand, models human cognition through symbolic reasoning combined with sub-symbolic activations, supporting sophisticated simulations of human cognitive processes like learning and memory.
+Its psychological plausibility makes it ideal for human-robot interaction scenarios, such as storytelling or social robotics, yet its heavy reliance on detailed knowledge engineering restricts flexibility in complex and dynamic environments.
+Finally, DIARC is explicitly designed for cognitive robotics with distributed modules [16, 17], affective processes, and advanced natural language capabilities suitable for dynamic human-robot interactions.
+However, it struggles with concurrent process management, lacks clear architectural layering, unified ontology, explainability, long-term episodic memory, explicit metacognition, and formalized safety mechanisms, thus limiting its robustness and adaptability.
+In summary, most architectures require extensive customization from both cognitive and control perspectives, leading to limited real-world applications and practical testing.
+II-B Multi-Agent Planning (MAP) and Communication
+Multi-Agent Planning (MAP) in Heterogeneous Multi-Robot (HMR) teams includes a variety of centralized and decentralized strategies, each presenting distinct trade-offs.
+Centralized planning achieves global optimality but encounters scalability challenges with larger teams, while decentralized methods improve scalability and adaptability but may result in suboptimal solutions [18], as highlighted by the exponential complexity identified by Brafman et al.
+[19].
+Consequently, hybrid MAP approaches that blend centralized hierarchical planning with decentralized execution [20], dynamically switch between plans [5], or integrate centralized conflict resolution with decentralized planning [21] have gained traction.
+Advanced implementations of these hybrid strategies often incorporate auction-based task allocation systems and distributed planners [22, 23].
+Regardless of the planning approach, effective intra-team communication remains integral for seamless coordination and execution within HMR teams.
+Such communication involves distinct protocols for human-robot and robot-robot interactions [24].
+A unified, human-centric communication approach within a common cognitive architecture is advocated as essential by Lemon et al.
 [25] and Natarajan et al.
-[5] support the need for a unified, human-centric approach to team communication.
-Rather than maintaining separate protocols for inter-robot and human-robot interactions, a common cognitive architecture that emulates human-like communication and reasoning offers a more cohesive solution, emulating human team dynamics.
-Early work by Clark et al.
-[21] proposed a framework for this process, which was later adapted by Klein et al.
-[22] to team coordination, introducing the notion of common ground.
-These studies highlight essential capabilities for robotic team members, including maintaining shared understanding, negotiating goals and plans, managing attention, and handling commands and requests.
-Subsequent pioneering research in dialogue processing systems [23, 24, 25] and grounding [26, 27, 28, 29] further developed these ideas, laying the groundwork for addressing communication challenges in human-robot teams.
-Results of this research and our own prior work on team organization [30] and natural language communication [7] guided the implementation of cognitive processes for team organization and communication in HARMONIC.
-III The HARMONIC Framework
-HARMONIC (Human-AI Robotic Team Member Operating with Natural Intelligence and Communication), shown in Fig.
-1, is a dual-control cognitive-robotic architecture that integrates strategic, cognitive-level decision-making with tactical, skill-level robot control.
-It builds upon and extends the concept of hybrid control systems and architectures, as discussed in a comprehensive review by Dennis et al.
 [7].
-It also represents an advancement over the type 2 integration approach employed in the DIARC framework [32, 33].
-Unlike DIARC, where the strategic layer is embedded as a subsystem within the tactical layer to enable concurrent and dynamic operations, HARMONIC introduces a more sophisticated integration of these components.
-The strategic and tactical architectural layers of HARMONIC are connected by a bidirectional interface for seamless communication.
-The strategic layer, built upon the OntoAgent cognitive architecture [34, 35, 7], encompasses modules for attention management, perception interpretation, and decision-making (see Section.
-IV).
-It employs both utility-based and analogical reasoning, enhanced by metacognitive abilities.
-This layer prioritizes strategic goals, manages plan agendas, and selects actions while continuously monitoring their execution.
-It also facilitates team-oriented operations, including communicating in natural language and explaining decisions.
-The tactical layer is responsible for robot control at a lower level, processing sensor inputs, and planning motor actions to execute high-level commands from the strategic layer.
-It employs controllers, algorithms, and models to translate abstract commands into concrete robot actions.
-For instance, a command PICK(key,at-position1) to ”pick up a key on the ground” initiates a series of precise operations, including object identification, position determination, trajectory computation, and actuation.
-This layer also manages reactive responses, such as collision avoidance, through dedicated controllers.
-In the tactical layer, HARMONIC employs Behavior Trees (BTs) [36], which provide modular and robust representations for skills and low-level plans (see Section.
-V).
-The architecture’s design aligns with Kahneman’s dual-system approach [37], with the strategic layer implementing System 2 (slow, deliberative reasoning) and the tactical layer implementing System 1 (fast, intuitive responses).
-This structure allows for dynamic scheduling and adaptation, enabling real-time adjustments to priorities and actions.
-It effectively handles computational delays, contingencies, safety concerns, and resource optimization through a combination of low-level planning and reactive algorithms.
-IV OntoAgent for Strategic Level Planning
-OntoAgent is a content-centric cognitive architecture incorporated in HARMONIC.
-OntoAgent is designed to support the development of social intelligent agents through computational cognitive modeling [35, 7, 34].
+Such architectures emulate human-like communication, reasoning, and team dynamics, building upon the notion of common ground introduced by Clark et al.
+[26] and further developed by Klein et al.
+[27].
+Subsequent advances in dialogue processing systems [28] and grounding techniques [29, 30], together with insights from prior work on team organization [31] and natural language communication [9], guide current implementations of cognitive processes of OntoAgent in HARMONIC.
+II-C The OntoAgent Cognitive Architecture
+OntoAgent is a content-centric cognitive architecture, designed to support the development of social intelligent agents through computational cognitive modeling [12, 9, 11].
 This approach emphasizes the need to acquire, maintain, and dynamically expand large-scale knowledge bases, which are essential for an agent’s perception interpretation, reasoning, and action.
@@
 It implements a graph database view of knowledge, allowing for complex relational queries and supporting the representation of inheritance hierarchies.
-OntoAgent places a strong emphasis on natural language understanding and meaning-based text generation.
+OntoAgent places a strong emphasis on natural language understanding (NLU) and meaning-based text generation.
 These capabilities rely on a semantic lexicon that links words and phrases with meanings grounded in a resident ontology.
 The language analyzer treats a large range of complex linguistic phenomena (lexical disambiguation, reference resolution, ellipsis reconstruction, new-word learning, etc.
-[7]) and generates ontologically-grounded interpretations of text meaning.
+[9, 12]) and generates ontologically-grounded interpretations of text meaning.
 The language generator, for its part, generates natural language utterances from ontologically grounded meaning representations that the agent creates as part of its reasoning about action.
@@
 It typically uses stored plans associated with goals but can also engage in reasoning from first principles when necessary.
-V Behavior Trees for Tactical Execution
+In the current work, OntoAgent is incorporated in the strategic layer of HARMONIC.
+II-D Behavior Trees
 Behavior Trees (BTs), originally developed for controlling non-player characters in video games, have found significant applications in robotics and AI [13].
 BTs are directed trees that provide a flexible and intuitive way to design control actions and define task-planning hierarchies for agents.
-The flexibility and modularity of BTs make them particularly suitable for complex, hierarchical task planning in robotics and AI applications, allowing for easy modification and scaling of agent behaviors.
-BTs in the tactical layer of HARMONIC enable effective reactive control and support the system’s safety and operational requirements in dynamic environments.
-The tactical layer also incorporates advanced control techniques such as whole-body compliant control [38, 36] with motion planning, path planning [39] etc.
-In addition to these, BTs can also be used for representing skills and low-level plans that can directly be shared between the robots[40, 41].
-BTs typically use a blackboard architecture to maintain and access frequently checked condition variables [42, 43].
-In the context of HARMONIC, a state manager keeps track of these variables on the tactical layer and allows for efficient querying and updating of the system’s state during operation based on the sensory inputs and action commands received from the strategic layer.
-Figure 2: (a) Root, control flow, and execution nodes in BTs. (b) BT design template for the tactical layer of robots running the HARMONIC architecture. (c) Sample BT on UGV.
-BTs for HARMONIC robots follow a design template prioritizing safety and needs [40], as shown in Fig.
-2.
-Collision avoidance subtrees are placed leftmost, followed by robot needs subtrees, then action command subtrees triggered by the strategic layer.
-Fallback subtrees on the far right provide actions when the robot is idle or other conditions aren’t met, such as random walks or waiting at the base station.
-This structure leverages the left-to-right execution of BTs to ensure proper prioritization of tasks.
-The translation of the action commands to the appropriate state flags and control variables is done through the action data communication APIs, which modify the placeholder variables in the State Manager on the tactical layer.
-VI Development of Distributed Plans
-In this section, we show how robots implemented in HARMONIC carry out a search task.
-They decide to act based on reasoning carried out in the strategic layer.
-Their plans (which are instances of ontological scripts) are divided into steps that are communicated to the tactical layer by means of action commands.
-VI-A High-level planning
+The structure of a BT includes control nodes (selectors, sequencers, and parallel nodes) that manage execution flow among child and execution nodes (action and condition nodes) that form the leaf nodes performing specific tasks.
+Control nodes process “ticks” at a set frequency, and conventionally execute from left to right, depending on the control node involved.
+This modular design enables easy modification and scaling of behaviors, making BTs particularly suitable for complex hierarchical task planning.
+In the current HARMONIC system, robot-specific BTs are integrated at the tactical layer to coordinate robotic control.
+III The HARMONIC Framework
+HARMONIC (Human-AI Robotic Team Member Operating with Natural Intelligence and Communication), shown in Fig.
+1, is a dual-control cognitive-robotic architecture that integrates strategic, cognitive-level decision-making with tactical, skill-level robot control.
+It builds upon and extends the concept of hybrid control systems and architectures, as discussed in a comprehensive review by Dennis et al.
+[32].
+It also represents an advancement over the type 2 integration approach employed in DIARC [16, 17].
+Unlike DIARC, where the strategic layer is embedded as a subsystem within the tactical layer to enable concurrent and dynamic operations, HARMONIC introduces a more sophisticated integration of these components.
+The strategic and tactical architectural layers of HARMONIC are connected by a bidirectional interface for seamless communication and data transfer.
+The strategic layer implements the OntoAgent cognitive architecture [11, 12, 9], encompasses modules for attention management, perception interpretation, and decision-making (see Sec.
+II-C).
+It employs both utility-based and analogical reasoning, enhanced by metacognitive abilities.
+This layer prioritizes strategic goals, manages plan agendas, and selects actions while continuously monitoring their execution.
+It also facilitates team-oriented operations, including communicating in natural language and explaining decisions.
+The tactical layer is responsible for attention management, lower-level robot planning and control, sensor input processing, and executing motor actions based on high-level commands received from the strategic layer.
+The attention manager processes perception data, translating and prioritizing sensory information for use by the strategic component.
+Additionally, the tactical layer employs specialized controllers, algorithms, and models to convert abstract commands from OntoAgent in the strategic layer into precise, actionable robot operations.
+For example, a command such as PICK(key, at-position1), from the OntoAgent to ”pick up a key on the ground,” triggers a well-defined sequence of tasks and actions including object identification, positional assessment, trajectory computation, and actuation, in the tactical layer.
+Furthermore, the tactical layer manages reactive behaviors, including collision avoidance, through dedicated controllers.
+The tactical layer of HARMONIC is built around the Behavior Trees (BTs) [13] which enable effective reactive control and support the system’s safety and operational requirements in dynamic environments.
+The tactical layer may also include advanced control policies for whole-body compliant control [33, 13] with motion planning, path planning [34] etc.
+In addition to these, BTs can also be used for representing skills and low-level plans that can also be shared between the robots [35, 36].
+In the context of HARMONIC, state manager (a blackboard) keeps track of condition and state variables on the tactical layer [37].
+As shown in Fig.
+2, this allows for efficient querying and updating of the system’s state during operation based on the sensory inputs and action commands received from the strategic layer.
+Sensory inputs are continuously written into the State Manager (SM), where the attention module filters and packages this perception data for the OntoAgent in the strategic layer.
+The high-level action commands sent by the OntoAgent through Interface APIs are unpacked for updating the state variables in the SM.
+While sensor data and physical state tracking operate in real time, the strategic layer runs asynchronously.
+Real-time control is achieved through the data available in the SM that is accessed by the LiveBT in parallel, which chooses policies and actions stored in the Action Manager.
+This architecture enables HARMONIC to maintain real-time behavior while supporting higher-level planning capabilities.
+Figure 2: Data flow and control diagram showing the connections between the BTs in the tactical layer and the OntoAgent in the strategic layer in the HARMONIC System.
+Action commands from the strategic layer are translated into corresponding state and control variables via action data communication APIs.
+These APIs update the placeholder variables within the State Manager, ensuring accurate and timely control of the tactical layer operations.
+The architecture’s design aligns with Kahneman’s dual-system approach [38], with the OntoAgent in the strategic layer implementing System 2 (slow, deliberative reasoning) and the BTs in the tactical layer implementing System 1 (fast, intuitive responses).
+This structure allows for dynamic scheduling and adaptation, enabling real-time adjustments to priorities and actions.
+It effectively handles computational delays, contingencies, safety concerns, and resource optimization through a combination of low-level planning and reactive algorithms.
+IV Task Design
+We evaluated the HARMONIC framework’s teaming strategy using a simulated search task set in an apartment environment, as illustrated in Fig.
+4.6.
+The scenario involves an HMR team consisting of an unmanned ground vehicle (UGV), a drone, and a human team member.
+The objective is to find a set of lost keys.
+The human initiates the task by remotely communicating with the robotic team members.
+The simulation environment, developed using the Unity game engine, incorporates abstracted versions of robot behaviors.
+This approach allowed us to focus on key aspects of the HARMONIC framework’s performance while simplifying the implementation of complex robot actions that are not central to our current work.
+IV-A Heterogenous Robot Team Configuration
+The simulation evaluation involved two robots: an Unmanned Ground Vehicle (UGV) and a drone, both functionally and structurally heterogeneous, offering complementary capabilities and perspectives [39].
+Each robot is equipped with sensors for localization, collision detection, and object recognition, facilitating coordinated and efficient operations.
+The UGV’s design (see Fig.
+3.a) enables it to access ground-level areas and maneuver under obstacles and surfaces.
+The drone, on the other hand, effectively complements the UGV by performing aerial search operations.
+The drone’s (see Fig.
+3.b) vertical mobility and sensing capabilities allow it to scan environments extensively from above, navigating over obstacles and accessing areas difficult for the UGV.
+Since the robots communicate in natural language, their dialogue and actions are comprehensible to humans, and humans can intervene when necessary.
+Both robots operate decentralized instances of tactical and strategic modules, ensuring efficient and autonomous team planning and control.
+Currently, the UGV acts as the team leader at the task level, responsible for managing interactions with the human operator and coordinating with the drone.
+The leader assignment is presently randomized; however, future work aims to develop and evaluate heuristic methods for dynamic team hierarchies.
+This advancement will optimize leader selection based on contextual factors and the individual capabilities of each agent relevant to specific tasks.
+IV-B Development of Distributed Plans
+Plans are distributed between strategic and tactical components in both UGV and drone systems, with each robot acting on reasoning from the OntoAgent in the strategic layer.
+The high-level plans (ontological script instances) are divided into steps communicated to the tactical layer as action commands, which are executed through low-level control encoded in BTs.
+IV-B1 High-level planning
 The ontologies of agents in HARMONIC contain scripts for complex events that a particular agent can perform or understand other agents performing.
-Plans are as instances of scripts, with parameter values set for the situation at hand.
-Plans are instantiated by executing one or more of ontologically stored meta-plans, that is, plans that HARMONIC robots instantiate to guide the generation of plans in their operational domain.
-One such meta-plan is the collaborative activity plan, whose purpose is to help agents operating in teams organize themselves to accomplish a goal.
-The details of the collaborative activity plan vary based on the agent’s role in the team.
+Plans are instances of scripts with parameter values set for the situation at hand.
+These are instantiated by executing one or more of the ontologically stored meta-plans, that is, plans that HARMONIC robots instantiate to guide the generation of plans in their operational domain.
+One such meta-plan is the COLLABORATIVE-ACTIVITY plan, whose purpose is to help agents operating in teams organize themselves to accomplish a goal.
+The details of the COLLABORATIVE-ACTIVITY plan vary based on the agent’s role in the team.
 For example, the team (or task) leader’s plan is typically more complex than the plans of its subordinates since team leaders select the plan to accomplish the target goal, resolve preconditions, and instruct their subordinates.
-Subordinates’ plans, by contrast, include awaiting instructions from the leader and, passively absorbing any other related information that comes their way.
+Subordinates’ plans, by contrast, include awaiting instructions from the leader and passively absorbing any other related information that comes their way.
 The leader’s script for collaboration is shown below.
-Since this is a meta-plan, it can be applied to many specific plans and, starts with few details populated.
+Since this is a meta-plan, it can be applied to many specific plans and starts with few details populated.
 Part of its task is to fill in task-specific details.
 Select comments are provided for this human-readable formalism.
-For further information, please refer to [34, 35, 7].
+For further details on this formalism, please refer to [11, 12, 9].
 {codeblock}
@@
 In the example below, the selected plan is SEARCH-FOR-LOST-OBJECT.
-The details of the selected plan can be placed into the [RUN-PLAN] section of collaborative activity, and any preconditions that must be resolved can be added to the [PRECONDITIONS] section.
+The details of the selected plan can be placed into the [RUN-PLAN] section of COLLABORATIVE-ACTIVITY, and any preconditions that must be resolved can be added to the [PRECONDITIONS] section.
 An excerpt from the leader’s script for COLLABORATIVE-ACTIVITY is shown below.
@@
 This process results in the team adopting a shared collaborative plan.
-VII Evaluation
-VII-A Simulation Environment and Task Design
-We evaluated the HARMONIC framework’s teaming strategy using a simulated search and retrieval task set in an apartment environment, as illustrated in Figure 5.
-The scenario involves a HMR team consisting of an unmanned ground vehicle (UGV), a drone, and a human team member serving as the team leader.
-The objective is to find a set of lost keys.
-The human initiates the task by remotely communicating with the robotic team members.
-The simulated apartment contains zones with varied robot accessibility, necessitating coordinated efforts.
-We categorize these as: a) spaces accessible to both robots, b) elevated surfaces accessible to only the drones, and c) spaces below surfaces accessible only to the UGVs, as shown in Fig.
-3.
-Each zone comprises waypoints - strategically defined locations that are dynamically utilized by the strategic layer.
-For this evaluation, we manually designated these waypoints and zones.
-However, in the future this process will be automated using machine learning models to classify and map zones based on spatial characteristics and accessibility, which will enhance the system’s adaptability to new environments.
-Figure 3: Zones in the search environment
-The simulation environment, developed using the Unity game engine, incorporates abstracted versions of robot behaviors to reduce complexity.
-This approach allowed us to focus on key aspects of the HARMONIC framework’s performance while simplifying the implementation of complex robot actions that are not central to our current work.
-VII-B Heterogenous Robot Team Configuration
-The simulations involved two robots, an Unmanned Ground Vehicle (UGV) and a drone (Fig.
-4).
-These robots are both functionally and structurally heterogeneous [44], thus providing a diverse range of complementary perspectives and capabilities to the team.
-Both robots are equipped with a suite of sensors that enable them with localization, collision detection, and object recognition capabilities.
-Figure 4: (a) UGV for ground-level exploration. (b) Parrot drone conducting an aerial scan of objects.
-Each robot runs its own instances of the tactical and strategic modules, enabling decentralized planning and control within the team.
-For the current scenario, the UGV is the robot team leader at the task level, making it responsible for maintaining interactions with the human and the drone.
-Currently, the leader assignment is randomized.
-In future work, we intend to develop and evaluate heuristic methods for implementing dynamic team hierarchies.
-This approach will optimize leadership selection based on contextual factors and individual agent capabilities depending on the task.
-In the tactical layer of both robots, sensor data is encoded into data frames and periodically shared with the strategic layer.
-Additionally, each robot incorporates reactive control strategies for collision avoidance, that are embedded within behavior trees in their respective tactical layers.
-The UGV is a custom-designed robot featuring a manipulator mounted atop a multidirectional mobile platform, as illustrated in Fig.
-4.a.
-The robot’s design allows it to access all ground-level locations and maneuver under certain objects, enhancing its capability to perform intricate search and manipulation tasks in confined spaces.
-Figure 2.c presents an excerpt from the Behavior Tree (BT) running on the Unmanned Ground Vehicle (UGV).
-This BT is based on the template shown in Figure 2.a.
-Figure 5: (1) Chat transcript between Danny and the robots.
-(2) UGV’s (leader) Text Meaning Representations (TMRs) (3, 6) A sample Vision Meaning Representations (VMRs) of detected objects by the robots (5) Leader’s agenda.
-(4, 7) Complete thought transcripts of UGV and Drone.
-The drone, on the other hand, can effectively perform search operations by scanning through a wide range of environments with sensors to scan the objects below it, as shown in Fig.
-4.b.
-Its vertical mobility allows it to navigate over obstacles and access areas that may be challenging for the UGV (type-b zones).
-Since the robots communicate in natural language, their dialog and actions are comprehensible to humans, and humans can intervene when necessary.
-VII-C Collaborative Search for Lost Keys
-The simulation environment is connected to a chatroom interface through which the human (Danny) communicates with the robots.
-A user interface (see Fig.
-5) includes not only a chat window but additional widgets that display the robots’ interpretations of language (TMRs) and visual inputs(VMRs), natural language renderings of the robots’ thoughts, and their goal and plan agendas.
-This arrangement illustrates the transparency of the robots’ operations and provides insights into their cognitive processes.
-The discussion below makes reference to the labels in Fig.
-5  [^1] A video of the complete simulation can be found at https://youtu.be/LpuE0yN0SEI
-VII-C1 Task Initiation
-The scenario begins with Danny, initiating the task by sending the robots a message (M1).
+IV-B2 Low-level control
+Behavior Trees (BTs) in the tactical layer for both robots follow a standardized design template prioritizing safety and operational needs as shown in Fig.
+3.c.
+This design leverages left-to-right execution priority, ensuring the robot’s critical needs and priorities.
+Collision avoidance sub-trees are positioned leftmost, followed by robot-needs sub-trees, then action-command sub-trees.
+On the far right, fallback sub-trees handle behaviors like random exploration or returning to the base station.
+An excerpt from the UGV’s BT is shown in Fig.
+3.d.
+Figure 3: (a) UGV for ground-level exploration. (b) Parrot drone conducting aerial scan. (c) BT design template for the tactical layer of robots. (d) Sample BT on UGV.
+Figure 4: (1) Chat transcript between Danny (human) and the robots.
+(2) UGV’s (leader) sample Vision Meaning Representations (VMRs) of detected objects by the robots (3) Leader’s agenda.
+(4, 5) Complete thought transcripts of UGV and Drone.
+(6) Multiple instances from the simulated apartment environment showing the UGV, the drone, and the keys.
+(i) Robots at their base stations.
+(ii, iii) Robots searching the apartment.
+(iv) UGV finds the keys.
+V Evaluation Results
+The simulation environment features an integrated chat UI through which the human operator (Danny) interacts with the robots.
+As shown in Fig.
+4(1-5), the interface comprises multiple components: a central chat window, visualization widgets for the robots’ language interpretations and visual perceptions (VMRs), natural language representations of robot cognition, and structured displays of their goals and plans.
+This comprehensive design offers transparency into the robots’ decision-making processes and cognitive operations.
+The analysis below refers to specific labeled elements in Fig.
+4.
+Please view the accompanying video for a better understanding.
+Additional information is also available on the HARMONIC project page.
+[^1] https://sanjayovs.github.io/HARMONIC2025IROS/
+V-1 Task Initiation
+The scenario begins with Danny initiating the task by sending the robots a message (M1).
 This input triggers the team leader (UGV) to place a COLLABORATIVE-ACTIVITY on its agenda (Fig.
-5.5) and launch the SEARCH-FOR-LOST-OBJECT plan.
+4.3) and launch the SEARCH-FOR-LOST-OBJECT plan.
 Concurrently, the drone awaits instructions.
-VII-C2 Information Gathering
+V-2 Information Gathering
 The UGV proceeds to verify preconditions for the SEARCH-FOR-LOST-OBJECT plan (Fig.
-5.5).
+4.3).
 Although the object type is already known, information about its features and last-seen location would be useful.
 The UGV queries Danny for this information (M2/M4), and receives responses (M3/M5).
-The latter response prompts the system to prioritize the entry-way sub-zone (type-a zone) in the search sequence.
-VII-C3 Search Execution
-With preconditions met, the UGV instructs the drone to initiate the search (M6), and both robots begin exploring their assigned areas.
+The latter response prompts the system to prioritize the entry-way in the search sequence (ii).
+V-3 Search Execution
+With preconditions met, the UGV instructs the drone to initiate the search (M6), and both robots begin exploring their assigned areas as shown in (ii-iv).
 The individual tactical modules on the robots control the search process using a waypoint strategy, while the strategic (cognitive) module maintains awareness of area existence without directly guiding robot navigation.
 This approach allows for efficient local path planning while preserving high-level planning in the strategic layer.
-VII-C4 Communication and Coordination
-Throughout the search, robots report their findings to each other and the human (M7-9). When a robot fails to locate the keys in a searched area, it communicates this to its partner.
-VII-C5 Task Completion
-The VMRs widget (Fig.
-5.3 and Fig.
-5.6) displays the object detection results that the strategic layer processes from sensing frames communicated by the tactical layer.
+V-4 Communication and Coordination
+Throughout the search, robots report their findings to each other and the human (M7-8, ii-iv). When a robot fails to locate the keys in a searched area, it communicates this to its partner.
+V-5 Task Completion
+The VMRs widget for UGV (Fig.
+4.2) displays the object detection results that the strategic layer processes from sensing frames communicated by the tactical layer.
 During the search execution, the strategic module continuously analyzes the sensor data frames, grounding these VMRs against the instance of the KEY object stored in its episodic memory.
-When the features match, the search is halted by either of the robots (leader in this case), informs the team (M8), and reports to Danny (M9).
+When the keys are found (iv) – the features match, the search is halted by either of the robots (leader in this case), informs the team (M8), and reports to Danny (M9).
 Notably, the UGV uses different language constructs when communicating with the drone versus Danny in (M9), demonstrating the cognitive agent’s ability to generate context-appropriate language.
-VIII Conclusion and Future Work
+VI Conclusion and Future Work
 This work presents a step towards the development of cognitive robots capable of effectively collaborating in heterogeneous human-robot teams.
-The HARMONIC architecture, with its dual-layer approach that combines strategic cognitive processing and tactical control enables the embodiment of cognitive architectures enhancing robots to exhibit metacognitive abilities, including team task understanding, episodic memory utilization, and multimodal perception interpretation.
-By integrating natural language communication and reasoning capabilities, our approach facilitates complex team interactions, enhances explainability, and builds trust the key factors in human-robot collaboration.
-The glass-box nature of our system, unlike black-box models based on LLMs or VLAs, provides transparency and interpretability to modules whose operation and output should be explainable to humans.
+The HARMONIC framework, with its dual-layer architecture that combines strategic cognitive processing and tactical control, allows robots to exhibit metacognitive abilities, including team task understanding, episodic memory utilization, and multimodal perception interpretation.
+By integrating natural language communication and reasoning capabilities, our approach facilitates complex team interactions, enhances explainability, and builds trust, the key factors in human-robot collaboration.
 Our demonstration system of a simulated search task involving two robots and a human teammate illustrates the potential for HARMONIC to be applied to real-world applications.
-Future research will focus on expanding the range of tasks and scenarios in which HARMONIC agents can successfully operate; enhancing the robots’ learning capabilities; conducting real-world trials using physical robots to validate the framework’s effectiveness in diverse operational contexts; developing collaborative online planning strategies; and, enhancing real-time adaptability in dynamic environments.
+Future research will focus on expanding the range of tasks and scenarios in which HARMONIC agents can successfully operate; enhancing the robots’ learning capabilities; conducting real-world trials using physical robots to validate the framework’s effectiveness in diverse operational contexts; developing collaborative online planning strategies; and enhancing real-time adaptability in dynamic environments.
 ACKNOWLEDGMENT

### v3

@@
 Abstract
-This paper introduces HARMONIC, a cognitive-robotic architecture that integrates the OntoAgent cognitive framework with general-purpose robot control systems applied to human-robot teaming (HRT).
-We also present a cognitive strategy for robots that incorporates metacognition, natural language communication, and explainability capabilities required for collaborative partnerships in HRT.
-Through simulation experiments involving a joint search task performed by a heterogeneous team of a UGV, a drone, and a human operator, we demonstrate the system’s ability to coordinate actions between robots with heterogeneous capabilities, adapt to complex scenarios, and facilitate natural human-robot communication.
-Evaluation results show that robots using the OntoAgent architecture within the HARMONIC framework can reason about plans, goals, and team member attitudes while providing clear explanations for their decisions, which are essential prerequisites for realistic human-robot teaming.
+This paper describes HARMONIC, a cognitive-robotic architecture that integrates the OntoAgent cognitive framework with general-purpose robot control systems applied to human-robot teaming (HRT).
+HARMONIC incorporates metacognition, meaningful natural language communication, and explainability capabilities required for developing mutual trust in HRT.
+Through simulation experiments involving a joint search task performed by a heterogeneous team of two HARMONIC-based robots and a human operator, we demonstrate heterogeneous robots that coordinate their actions, adapt to complex scenarios, and engage in natural human-robot communication.
+Evaluation results show that HARMONIC-based robots can reason about plans, goals, and team member attitudes while providing clear explanations for their decisions, which are essential requirements for realistic human-robot teaming.
 I Introduction
-The proliferation of robots across various domains has given rise to a growing number of human-robot teams, where robots with various capabilities are employed by human operators to achieve complex mission goals [1].
-Current state-of-the-art robotic systems demonstrate remarkable tactical proficiency, from the sub-millimeter precision of surgical robots performing intricate surgical procedures [2] to the dynamic multi-terrain locomotion of quadrupedal [3] and humanoid platforms exhibiting sophisticated whole-body coordination [4].
-These systems also excel at integrating multi-modal sensory information through sensor fusion, simultaneously localizing themselves within complex environments, recognizing objects, tracking moving targets with increasing reliability, in adaptive policy learning and multi-robot planning and collaboration [5, 6].
+Current state-of-the-art robotic systems demonstrate remarkable tactical proficiency, from the sub-millimeter precision of surgical robots performing surgical procedures [2] to the dynamic multi-terrain locomotion of quadrupedal [3] and humanoid platforms exhibiting sophisticated whole-body coordination [4].
+These systems also excel at integrating multimodal sensory information through sensor fusion, simultaneously localizing themselves within complex environments, recognizing objects, tracking moving targets with increasing reliability, in adaptive policy learning and multi-robot planning and collaboration [5, 6].
 Despite these significant advancements in tactical execution, most state-of-the-art robots are limited to short-horizon planning and execution and lack the strategic capabilities necessary for long-horizon planning, reasoning, and true collaborative partnership with humans in complex decision-making scenarios [7].
-Importantly, effective human-robot teaming requires more than just tactical proficiency; to function as trusted partners within human-robot teams, they must exhibit a range of cognitive and metacognitive capabilities, including:
-  • developing and enhancing an understanding of team tasks and team organization;
-  • developing and enhancing specific capabilities, involving the responsibilities, preferences, and actions of self and other team members [8];
-  • maintaining and using an episodic memory of past activities, as when leveraging past experiences to make informed predictions about future needs and actions;
-  • interpreting language, visual perception, and haptic perception in terms of an underlying formal model of the world (ontology); [9]
-  • communicating with teammates using interpreted, meaningful natural language.
-Figure 1: An overview of the framework, showing the Strategic and Tactical components representing the high-level planning (System 2) and low-level execution (System 1), respectively.
-Long-horizon functioning is within the scope of the field of cognitive architectures where much progress has been made in supporting complex agent team interactions, managing attention, setting goals, and conducting sophisticated planning, often incorporating diverse multi-agent planning strategies.
-Still, while cognitive architectures address key limitations in current robotic systems, they do not traditionally concentrate on the dynamic world of embodied agents.
-This has resulted in a notable disconnect between cognitive architectures and practical robotic control systems.
-Bridging this divide leverages the strengths of both these areas of research and development.
-The HARMONIC framework presented in this paper directly addresses this challenge through the use of a dual-layer architecture that combines strategic cognitive capabilities with tactical robotic control.
+Importantly, effective human-robot teaming [1] requires more than just tactical proficiency; to function as trusted partners within human-robot teams, they must exhibit a range of cognitive and metacognitive capabilities, including:
+  • developing and refining an understanding of team tasks and organization;
+  • building specific capabilities related to the responsibilities, preferences, and actions of self and other team members [8];
+  • maintaining and using episodic memory to leverage past experiences for predicting future needs and actions;
+  • interpreting language, visual, and haptic perception through an underlying formal model of the world (ontology) [9];
+  • communicating with teammates using meaningful, interpreted natural language.
+Figure 1: An overview of the HARMONIC framework, showing the Strategic and Tactical components representing the high-level planning (System 2) and low-level execution (System 1), respectively.
+Long-horizon functioning is within the scope of the field of cognitive systems, where much progress has been made on the tasks of complex agent team interactions, managing attention, goal selection, and plan execution.
+Still, while cognitive systems address key limitations in current robotic systems, they do not traditionally concentrate on the dynamic world of embodied agents.
+This has resulted in a disconnect between cognitive systems and practical robotic control systems.
+Our work integrates and leverages the strengths of both these areas of research and development.
+The HARMONIC framework uses a dual-layer architecture that integrates strategic cognitive capabilities with tactical robotic control.
 This integration not only enables long-horizon planning but also enhances the scope of the robot’s reasoning capabilities, adds metacognitive awareness and social intelligence, which in turn supports human-level explanation that serves as a foundation for robots to function as trusted teammates [10], rather than mere tools, within human-robot teams.
-The cognitive architecture that we incorporate into HARMONIC as its strategic layer is OntoAgent [11, 12, 9].
-It is responsible for the aforementioned cognitive architecture capabilities, thus facilitating complex team interactions, attention management, goal-setting, sophisticated planning (including the incorporation of diverse multi-agent planning strategies), and addressing unexpected challenges in ways that humans can readily interpret.
-Additionally, OntoAgent integrates both verbal and visual communication, significantly enhancing skill transfer and knowledge sharing within human-robot teams, grounding, and supporting meaning-oriented natural language communication, enabling clear causal explanations for the robots’ actions and decisions.
-These capabilities greatly enhance transparency and explainability [10], establishing crucial foundations for trust in human-robot collaborations [9].
+HARMONIC incorporates the cognitive architecture OntoAgent [11, 12, 9] as a strategic layer.
+Within HARMONIC, OntoAgent is responsible for semantic interpretation of sensory perception, grounding, attention management, goal-setting, sophisticated planning (supporting diverse multi-agent planning strategies), and addressing unexpected challenges in ways that humans can readily interpret.
+Additionally, OntoAgent supports meaning-oriented natural language communication, enabling clear causal explanations for the robots’ actions and decisions [10].
+This capability significantly enhances skill transfer and knowledge sharing within human-robot teams.
+The use of the OntoAgent framework greatly enhances transparency of decision-making in HARMONIC-based robots, establishing crucial foundations for trust in human-robot collaboration [9].
 We present our approach to human-robot teaming using the example of a team comprised of two robots and a human executing a joint task.
-We demonstrate the robots’ knowledge of team organization as well as, their ability to (a) dynamically negotiate tasks and allocate responsibilities, (b) select and modify plans, and (c) adapt language communication to both the situation and the needs of their interlocutors.
-We implement these robotic capabilities using the novel HARMONIC architecture that operates in parallel at two levels – the strategic cognitive level, which supports operations requiring reasoning, and the tactical control level, which supports skill-based, ”automatic,” reflexive operation of robots.
-Specifically, the key contributions of this paper are:
-  • We introduce HARMONIC, a cognitive-robotic architecture, which enables a flexible integration of OntoAgent [11], a cognitive architecture with low-level robot planning and control using Behavior Trees [13].
+We demonstrate the robots’ knowledge of team organization as well as their ability to (a) dynamically negotiate tasks and allocate responsibilities, (b) select and modify plans, and (c) adapt language communication to both the situation and the needs of their interlocutors.
+We implement these robotic capabilities using the novel HARMONIC architecture that operates in parallel at two levels – the strategic cognitive level, which supports operations requiring reasoning, and the tactical control level, which supports skill-based, “automatic,” reflexive operation of robots.
+Specifically, the key contributions of this paper are as follows:
+  • We introduce HARMONIC, a cognitive-robotic architecture, which enables a flexible integration of OntoAgent [11], a cognitive architecture with robot planning and control using Behavior Trees [13].
   • A cognitive strategy for multi-robot planning and execution with metacognition, communication, and explainability.
 This includes the robots’ natural language understanding and generation, their reasoning about plans, goals, and attitudes, and their ability to explain the reasons for their own and others’ actions.
-  • Results of simulation experiments on a search task carried out by a robotic team, which showcases the human-robot team’s interactions and ability to handle complex, real-world scenarios.
+  • Results of simulation experiments on a search task carried out by a robotic team, which showcases the human-robot team’s interactions and ability to handle complex, scenarios.
 Section II briefly describes the state of the art and background.
 Section III introduces the HARMONIC framework and the implementation of its strategic and tactical levels.
-Section IV-B describes the task, the team configuration, and the development of distributed plans, and Section IV presents the evaluation.
+Section IV describes the task, the team configuration, and the development of distributed plans, and Section V presents the evaluation.
 II Background and Related Work
 II-A Cognitive-Robotic Architectures
-There are relatively few cognitive architectures explicitly designed or effectively adapted for robotic applications due to the significant challenges involved in combining cognitive-level decision-making with motor control.
+There are relatively few cognitive architectures explicitly designed or effectively adapted for robotic applications.
 Among these architectures, Soar [14], ACT-R [15], and DIARC [16] are prominent, each contributing distinct methods and philosophies toward cognitive autonomy in robotics.
@@
 Its psychological plausibility makes it ideal for human-robot interaction scenarios, such as storytelling or social robotics, yet its heavy reliance on detailed knowledge engineering restricts flexibility in complex and dynamic environments.
-Finally, DIARC is explicitly designed for cognitive robotics with distributed modules [16, 17], affective processes, and advanced natural language capabilities suitable for dynamic human-robot interactions.
-However, it struggles with concurrent process management, lacks clear architectural layering, unified ontology, explainability, long-term episodic memory, explicit metacognition, and formalized safety mechanisms, thus limiting its robustness and adaptability.
+DIARC is explicitly designed for cognitive robotics with distributed modules [16, 17], affective processes, and natural language capabilities suitable for dynamic human-robot interactions in narrow domains.
+However, it struggles with concurrent process management and lacks clear architectural layering, unified ontology, explainability, long-term episodic memory, explicit metacognition, and formalized safety mechanisms, thus limiting its robustness and adaptability.
 In summary, most architectures require extensive customization from both cognitive and control perspectives, leading to limited real-world applications and practical testing.
 II-B Multi-Agent Planning (MAP) and Communication
-Multi-Agent Planning (MAP) in Heterogeneous Multi-Robot (HMR) teams includes a variety of centralized and decentralized strategies, each presenting distinct trade-offs.
-Centralized planning achieves global optimality but encounters scalability challenges with larger teams, while decentralized methods improve scalability and adaptability but may result in suboptimal solutions [18], as highlighted by the exponential complexity identified by Brafman et al.
-[19].
-Consequently, hybrid MAP approaches that blend centralized hierarchical planning with decentralized execution [20], dynamically switch between plans [5], or integrate centralized conflict resolution with decentralized planning [21] have gained traction.
-Advanced implementations of these hybrid strategies often incorporate auction-based task allocation systems and distributed planners [22, 23].
-Regardless of the planning approach, effective intra-team communication remains integral for seamless coordination and execution within HMR teams.
-Such communication involves distinct protocols for human-robot and robot-robot interactions [24].
-A unified, human-centric communication approach within a common cognitive architecture is advocated as essential by Lemon et al.
+Multi-Agent Planning (MAP) in Heterogeneous Multi-Robot (HMR) teams includes a variety of centralized, decentralized, and hybrid strategies[20, 5, 21], each presenting distinct trade-offs.
+None of the current MAP approaches incorporates sufficient means of intra-team communication, which is an essential requirement for robust HMR teaming.
+Such communication must involve distinct protocols for human-robot and robot-robot interactions [24].
+A unified, human-centric communication approach within a common cognitive architecture is advocated by Lemon et al.
 [25] and Natarajan et al.
 [7].
-Such architectures emulate human-like communication, reasoning, and team dynamics, building upon the notion of common ground introduced by Clark et al.
+HARMONIC, with OntoAgent as its cognitive strategic layer, implements human-oriented communication, reasoning, and team dynamics.
+It also incorporates a variety of these scientific advances, such as the notion of common ground introduced by Clark et al.
 [26] and further developed by Klein et al.
-[27].
-Subsequent advances in dialogue processing systems [28] and grounding techniques [29, 30], together with insights from prior work on team organization [31] and natural language communication [9], guide current implementations of cognitive processes of OntoAgent in HARMONIC.
+[27], the latest in dialogue processing systems [28], grounding techniques [29, 30], work on team organization [31] and natural language communication [9].
 II-C The OntoAgent Cognitive Architecture
-OntoAgent is a content-centric cognitive architecture, designed to support the development of social intelligent agents through computational cognitive modeling [12, 9, 11].
+OntoAgent is a cognitive architecture, designed to support the development of social intelligent agents supporting content-centric computational cognitive modeling [12, 9, 11].
 This approach emphasizes the need to acquire, maintain, and dynamically expand large-scale knowledge bases, which are essential for an agent’s perception interpretation, reasoning, and action.
@@
 The architecture is built on a service-based infrastructure, with key services including perception interpretation, attention management, reasoning, and action rendering.
-OntoAgent can integrate both native services and imported external capabilities, such as robotic vision or test-to-speech modules.
+OntoAgent can integrate both native services and imported external capabilities, such as robotic vision or text-to-speech modules.
 This flexibility allows it to support multiple perception channels, including language understanding, visual perception, and simulated interoception (Chapter 8 of [12]).
@@
 In the current work, OntoAgent is incorporated in the strategic layer of HARMONIC.
-II-D Behavior Trees
-Behavior Trees (BTs), originally developed for controlling non-player characters in video games, have found significant applications in robotics and AI [13].
-BTs are directed trees that provide a flexible and intuitive way to design control actions and define task-planning hierarchies for agents.
-The structure of a BT includes control nodes (selectors, sequencers, and parallel nodes) that manage execution flow among child and execution nodes (action and condition nodes) that form the leaf nodes performing specific tasks.
-Control nodes process “ticks” at a set frequency, and conventionally execute from left to right, depending on the control node involved.
-This modular design enables easy modification and scaling of behaviors, making BTs particularly suitable for complex hierarchical task planning.
-In the current HARMONIC system, robot-specific BTs are integrated at the tactical layer to coordinate robotic control.
 III The HARMONIC Framework
@@
 [32].
-It also represents an advancement over the type 2 integration approach employed in DIARC [16, 17].
+It also represents an advance over the type 2 integration approach employed in DIARC [16, 17].
 Unlike DIARC, where the strategic layer is embedded as a subsystem within the tactical layer to enable concurrent and dynamic operations, HARMONIC introduces a more sophisticated integration of these components.
@@
 It also facilitates team-oriented operations, including communicating in natural language and explaining decisions.
-The tactical layer is responsible for attention management, lower-level robot planning and control, sensor input processing, and executing motor actions based on high-level commands received from the strategic layer.
+When the strategic layer receives perception inputs, including speech, vision, and robot state information, as data frames from the tactical layer, it interprets them in the context of its active goals and situation model.
+These inputs are transformed into ontologically grounded meaning representations, enabling unified reasoning across modalities.
+This interpretation supports downstream processes such as attention management and strategic decision-making by allowing the agent to evaluate the relevance of new information, update its understanding of the environment, monitor progress on current plans, and determine whether adjustments are necessary.
+Based on this reasoning, the OntoAgent in the strategic layer may issue updated action commands or generate natural language responses to teammates.
+The tactical layer, on the other hand, is responsible for reflexive attention management, robot action planning and control, sensor input processing, and executing motor actions based on high-level commands received from the strategic layer.
 The attention manager processes perception data, translating and prioritizing sensory information for use by the strategic component.
 Additionally, the tactical layer employs specialized controllers, algorithms, and models to convert abstract commands from OntoAgent in the strategic layer into precise, actionable robot operations.
-For example, a command such as PICK(key, at-position1), from the OntoAgent to ”pick up a key on the ground,” triggers a well-defined sequence of tasks and actions including object identification, positional assessment, trajectory computation, and actuation, in the tactical layer.
+For example, a command such as PICK(key, at-position1), from the OntoAgent to “pick up a key on the ground,” triggers a well-defined sequence of tasks and actions, including object identification, positional assessment, trajectory computation, and actuation, in the tactical layer.
 Furthermore, the tactical layer manages reactive behaviors, including collision avoidance, through dedicated controllers.
-The tactical layer of HARMONIC is built around the Behavior Trees (BTs) [13] which enable effective reactive control and support the system’s safety and operational requirements in dynamic environments.
-The tactical layer may also include advanced control policies for whole-body compliant control [33, 13] with motion planning, path planning [34] etc.
-In addition to these, BTs can also be used for representing skills and low-level plans that can also be shared between the robots [35, 36].
-In the context of HARMONIC, state manager (a blackboard) keeps track of condition and state variables on the tactical layer [37].
+The tactical layer of HARMONIC is built around the Behavior Trees (BTs) [13], which enable effective reactive control and support the system’s safety and operational requirements in dynamic environments.
+The tactical layer may also include advanced control policies for whole-body compliant control [33, 13] with motion planning, path planning [34], etc.
+BTs can also be used for representing skills and low-level plans that can also be shared between the robots [35, 36].
+In the context of HARMONIC, the state manager module (a blackboard) keeps track of condition and state variables on the tactical layer [37].
 As shown in Fig.
@@
 Sensory inputs are continuously written into the State Manager (SM), where the attention module filters and packages this perception data for the OntoAgent in the strategic layer.
-The high-level action commands sent by the OntoAgent through Interface APIs are unpacked for updating the state variables in the SM.
+Simultaneously, the high-level action commands sent by the OntoAgent through the Interface APIs are also unpacked for updating the variables in the SM.
 While sensor data and physical state tracking operate in real time, the strategic layer runs asynchronously.
-Real-time control is achieved through the data available in the SM that is accessed by the LiveBT in parallel, which chooses policies and actions stored in the Action Manager.
+Real-time control is achieved through the data available in the SM, which is accessed in parallel by the LiveBT (a BT instance), a module that chooses policies and actions stored in the Action Manager libraries.
 This architecture enables HARMONIC to maintain real-time behavior while supporting higher-level planning capabilities.
 Figure 2: Data flow and control diagram showing the connections between the BTs in the tactical layer and the OntoAgent in the strategic layer in the HARMONIC System.
-Action commands from the strategic layer are translated into corresponding state and control variables via action data communication APIs.
 These APIs update the placeholder variables within the State Manager, ensuring accurate and timely control of the tactical layer operations.
-The architecture’s design aligns with Kahneman’s dual-system approach [38], with the OntoAgent in the strategic layer implementing System 2 (slow, deliberative reasoning) and the BTs in the tactical layer implementing System 1 (fast, intuitive responses).
-This structure allows for dynamic scheduling and adaptation, enabling real-time adjustments to priorities and actions.
-It effectively handles computational delays, contingencies, safety concerns, and resource optimization through a combination of low-level planning and reactive algorithms.
+The HARMONIC architecture aligns with Kahneman’s dual-system approach [38], with the OntoAgent in the strategic layer implementing System 2 (slow, deliberative reasoning) and the BTs in the tactical layer implementing System 1 (fast, intuitive responses).
+This structure allows for integrating reflexive and deliberative functioning, flexible scheduling and adaptation, which in turn enables real-time prioritization of goals and actions.
+As a result, HARMONIC effectively handles computational delays, contingencies, safety concerns, and resource optimization through a combination of low-level planning and reactive algorithms.
 IV Task Design
@@
 The objective is to find a set of lost keys.
-The human initiates the task by remotely communicating with the robotic team members.
-The simulation environment, developed using the Unity game engine, incorporates abstracted versions of robot behaviors.
+The human initiates the task by remotely communicating with the robotic team members in English.
+The simulation environment, developed in Unity, incorporates abstracted versions of robot behaviors.
 This approach allowed us to focus on key aspects of the HARMONIC framework’s performance while simplifying the implementation of complex robot actions that are not central to our current work.
 IV-A Heterogenous Robot Team Configuration
-The simulation evaluation involved two robots: an Unmanned Ground Vehicle (UGV) and a drone, both functionally and structurally heterogeneous, offering complementary capabilities and perspectives [39].
+The simulation evaluation involved two functionally and structurally heterogeneous robots – a UGV and a drone – offering complementary capabilities and perspectives [39].
 Each robot is equipped with sensors for localization, collision detection, and object recognition, facilitating coordinated and efficient operations.
-The UGV’s design (see Fig.
-3.a) enables it to access ground-level areas and maneuver under obstacles and surfaces.
-The drone, on the other hand, effectively complements the UGV by performing aerial search operations.
+The UGV’s (see Fig.
+3.a) can access ground-level areas and maneuver under obstacles and surfaces.
+The drone, on the other hand, complements the UGV by performing aerial search operations.
 The drone’s (see Fig.
-3.b) vertical mobility and sensing capabilities allow it to scan environments extensively from above, navigating over obstacles and accessing areas difficult for the UGV.
+3.b) vertical mobility and sensing capabilities allow it to scan environments extensively from above, navigating over obstacles and accessing areas difficult for the UGV to access.
 Since the robots communicate in natural language, their dialogue and actions are comprehensible to humans, and humans can intervene when necessary.
-Both robots operate decentralized instances of tactical and strategic modules, ensuring efficient and autonomous team planning and control.
-Currently, the UGV acts as the team leader at the task level, responsible for managing interactions with the human operator and coordinating with the drone.
-The leader assignment is presently randomized; however, future work aims to develop and evaluate heuristic methods for dynamic team hierarchies.
+Both robots operate independent instances of HARMONIC, with different content parts (for example, there is no presumption that robots have identical ontological world models).
+For the experiment, the UGV acts as the team leader at the task level, responsible for managing interactions with the human operator and coordinating with the drone.
+The team leader assignment is presently arbitrary; however, future work aims to develop and evaluate heuristic methods for dynamic team hierarchies.
 This advancement will optimize leader selection based on contextual factors and the individual capabilities of each agent relevant to specific tasks.
 IV-B Development of Distributed Plans
-Plans are distributed between strategic and tactical components in both UGV and drone systems, with each robot acting on reasoning from the OntoAgent in the strategic layer.
-The high-level plans (ontological script instances) are divided into steps communicated to the tactical layer as action commands, which are executed through low-level control encoded in BTs.
+Plans are distributed between strategic and tactical components in both the UGV and the drone.
+The high-level plans (ontological script instances) are divided into steps communicated to the tactical layer as action commands, as described earlier.
 IV-B1 High-level planning
@@
 Plans are instances of scripts with parameter values set for the situation at hand.
-These are instantiated by executing one or more of the ontologically stored meta-plans, that is, plans that HARMONIC robots instantiate to guide the generation of plans in their operational domain.
-One such meta-plan is the COLLABORATIVE-ACTIVITY plan, whose purpose is to help agents operating in teams organize themselves to accomplish a goal.
-The details of the COLLABORATIVE-ACTIVITY plan vary based on the agent’s role in the team.
-For example, the team (or task) leader’s plan is typically more complex than the plans of its subordinates since team leaders select the plan to accomplish the target goal, resolve preconditions, and instruct their subordinates.
-Subordinates’ plans, by contrast, include awaiting instructions from the leader and passively absorbing any other related information that comes their way.
+These are instantiated by executing one or more of the ontologically stored meta-scripts, that is, scripts that HARMONIC robots instantiate to guide the generation of plans from scripts in their operational domain.
+One such meta-script is the COLLABORATIVE-ACTIVITY script, whose purpose is to help agents operating in teams organize themselves to accomplish a goal.
+The details of the COLLABORATIVE-ACTIVITY script vary based on the agent’s role in the team.
+For example, the team (or task) leader’s script is typically more complex than the scripts of its subordinates since team leaders must select a domain script to instantiate into the plan to accomplish the target goal, resolve preconditions, and instruct their subordinates.
+Subordinates’ scripts, by contrast, include awaiting instructions from the leader and passively absorbing any other related information that comes their way.
 The leader’s script for collaboration is shown below.
-Since this is a meta-plan, it can be applied to many specific plans and starts with few details populated.
-Part of its task is to fill in task-specific details.
+Since this is a meta-script, it can be applied to many specific scripts and starts with few parameter values populated.
+Part of its task is to fill in task-specific parameter values.
 Select comments are provided for this human-readable formalism.
@@
 AWAIT \$.HAS-COLLABORATIVE-PLAN ISA @EVENT [RUN-PLAN] // The subordinate has no plan yet.
-As the leader begins to process the meta-plan, it will select the domain plan to pursue.
-In the example below, the selected plan is SEARCH-FOR-LOST-OBJECT.
-The details of the selected plan can be placed into the [RUN-PLAN] section of COLLABORATIVE-ACTIVITY, and any preconditions that must be resolved can be added to the [PRECONDITIONS] section.
+Having instantiated an instance of COLLABORATIVE-ACTIVITY as a meta-plan, the leader will select the domain script to pursue and instantiate as a domain plan.
+In the example below, the selected script is SEARCH-FOR-LOST-OBJECT.
+The details of the plan instantiated from the selected script can be placed into the [RUN-PLAN] section of COLLABORATIVE-ACTIVITY, and any preconditions that must be resolved can be added to the [PRECONDITIONS] section.
 An excerpt from the leader’s script for COLLABORATIVE-ACTIVITY is shown below.
@@
 Once the leader has worked through the preconditions, it can move to SUGGEST-PLAN, in which it will share the domain plan with subordinates through natural language dialog.
-These agents will receive this information and adopt their assignment(s) within the suggested plan.
-This process results in the team adopting a shared collaborative plan.
+These subordinates will receive this information and adopt their assignment(s) within the suggested plan, resulting in the team adopting a shared collaborative plan.
 IV-B2 Low-level control
-Behavior Trees (BTs) in the tactical layer for both robots follow a standardized design template prioritizing safety and operational needs as shown in Fig.
+For low-level planning and control on the tactical layer, we use Behavior Trees (BTs).
+BTs provide a flexible and intuitive way to design control actions and define task-planning hierarchies for robots [13].
+Their modular design enables easy modification and scaling of behaviors, making BTs particularly suitable for complex hierarchical task planning, in the tactical layer.
+BTs for both robots follow a standardized design template prioritizing safety and operational needs as shown in Fig.
 3.c.
-This design leverages left-to-right execution priority, ensuring the robot’s critical needs and priorities.
-Collision avoidance sub-trees are positioned leftmost, followed by robot-needs sub-trees, then action-command sub-trees.
-On the far right, fallback sub-trees handle behaviors like random exploration or returning to the base station.
+This design leverages the conventional left-to-right execution priority, ensuring the robot’s critical needs and priorities.
+Collision avoidance subtrees are positioned leftmost, followed by robot-needs sub-trees, then action-command sub-trees.
+On the far right, fallback sub-trees handle behaviors like random exploration.
 An excerpt from the UGV’s BT is shown in Fig.
@@
 4(1-5), the interface comprises multiple components: a central chat window, visualization widgets for the robots’ language interpretations and visual perceptions (VMRs), natural language representations of robot cognition, and structured displays of their goals and plans.
-This comprehensive design offers transparency into the robots’ decision-making processes and cognitive operations.
+This design offers transparency into the robots’ decision-making processes and cognitive operations.
 The analysis below refers to specific labeled elements in Fig.
@@
 Please view the accompanying video for a better understanding.
-Additional information is also available on the HARMONIC project page.
-[^1] https://sanjayovs.github.io/HARMONIC2025IROS/
 V-1 Task Initiation
@@
 The UGV queries Danny for this information (M2/M4), and receives responses (M3/M5).
-The latter response prompts the system to prioritize the entry-way in the search sequence (ii).
+The latter response prompts the system to prioritize entry-way in the search sequence (ii).
 V-3 Search Execution
@@
 By integrating natural language communication and reasoning capabilities, our approach facilitates complex team interactions, enhances explainability, and builds trust, the key factors in human-robot collaboration.
-Our demonstration system of a simulated search task involving two robots and a human teammate illustrates the potential for HARMONIC to be applied to real-world applications.
-Future research will focus on expanding the range of tasks and scenarios in which HARMONIC agents can successfully operate; enhancing the robots’ learning capabilities; conducting real-world trials using physical robots to validate the framework’s effectiveness in diverse operational contexts; developing collaborative online planning strategies; and enhancing real-time adaptability in dynamic environments.
+Our demonstration system of a simulated search task involving two robots and a human teammate illustrates the potential for HARMONIC to be applied to MRS problems.
+Future research will focus on expanding the range of tasks and scenarios in which HARMONIC agents can successfully operate; enhancing the robots’ learning capabilities; conducting real-world trials using physical robots to validate the framework’s effectiveness in diverse operational contexts; and developing collaborative online planning strategies in dynamic environments;
 ACKNOWLEDGMENT
-This research was supported in part by grant ##N00014-23-1-2060 from the U.S.
-Office of Naval Research.
-Any opinions or findings expressed in this material are those of the authors and do not necessarily reflect the views of the Office of Naval Research.
+This work was supported in part by ONR Grant #N00014-23-1-2060. The views expressed are those of the authors and do not necessarily reflect those of the Office of Naval Research.